OpenStudy (anonymous):

A rectangle has a length of 18 cm and a perimeter of 68 cm. Which equation could be used to find the width of the rectangle? A. w + 18 = 68 B. 18w = 68 C. 2w + 18 = 68 D. 2w + 36 = 68

jigglypuff314 (jigglypuff314):

Perimeter of Rectangle = 2w +2L plug in what you know 68 = 2w + 2(18) and simplify :)

OpenStudy (anonymous):

2w +2L = perimeter so.... 2w + 36 = 68

OpenStudy (anonymous):

iz back

OpenStudy (anonymous):

and jiggly its peldo:D

OpenStudy (anonymous):

u get metal

OpenStudy (cakt07):

A?

OpenStudy (cakt07):

no i think B

OpenStudy (anonymous):

.

OpenStudy (mathmale):

Both jigglypuff and skinny are right on target, @cakt07 and @ijspeifioef. Why not build on their work, showing how you'd solve for w?